You are here: CLASSE Wiki>ACC/Bunch Web>BSM23WLog>BSM23WLog_2006 (21 May 2007, EugeneTanke)Edit Attach
Tags

BSM23W Vertical Beamsize Monitor Electronic Logbook 2006

23 Jul 2024 - 02:39

January 2006

January 20, 2006 (E. Tanke): CHESS X-ray BSM

  • This week cables have been pulled to CHESS hutch B1 for CHESS X-ray BSM experiments. A database node has been added to talk to the DSP at this location: BSM PROC XXX 3, where XXX=ADR, PKT and DAT.
  • With a test signal (pulse length ~ 4ns) supplied by a LeCroy pulse generator connected to the diode board (no diode mounted yet, instead a 50 Ohm resister to ground and a 1.2 pF to the input were mounted), the following response was found:
    Xray BSM response
  • For future reference, see the CHESS X-Ray BSM TWIKI page

February 2006

February 16 , 2006 (M.Palmer, E.Tanke) BSM23W beam size measurements

  • Measurements with bsm_v1 (averaging routine), PMT HV=420V, Tglob=1850, e-= 177 mA, e+= 125 mA, bunches selected: 6 by 6.
    • 10:59 Measurement stored in /home/tanke/cesr/ts101/bsm23w/test_300avg/2006_02_16/1059, 100 loops
    • 11:11 Measurement stored in /home/tanke/cesr/ts101/bsm23w/test_300avg/2006_02_16/1111, 300 loops
    • Fits made for channels 7 to 16 (filenames contain _7t16) and for channels 11 to 15 ( _11t15)

February 27 , 2006 (R.Eshelman, E.Tanke) BSM23W beam size measurements

  • Measurements with bsm_v1 (averaging routine), PMT HV=400V, Tglob=1820, 100 loops, bunches selected: 6 by 6. Two measurements were made: one at the beginning of a run (19:19) and one at the end (23:00). Mirror positions: H=Touchy 1=530, V=Touchy 2=541
    • 19:19 Results stored in /home/tanke/cesr/ts101/bsm23w/test_300avg/2006_02_27/1919, e- = 227 mA
    • 23:00 Results stored in /home/tanke/cesr/ts101/bsm23w/test_300avg/2006_02_27/2300, e- = 187 mA
    • Analyzed results can be found in: ms_2006_02_27

March 2006

March 7, 2006 Installation of 23W Horizontal Periscope (MAP/RLH)

  • Plan:
    • Set up laser optics and place marker flags internal to light box
    • Replace horizontal periscope
    • Realign to laser flags
  • Installation ok, but laser was removed by tunnel crew thus preventing us recovering alignment
  • Will need to do full alignment 3/7-3/8.

March 8, 2006 Attempt to Realign e- Synchrotron Light Monitor Optics (MAP/EPT/RME/RNC/LMW)

  • Full description with pictures can be found here ( takes time to download this page )

March 15, 2006 Ecloud machine study (J.Codner, D.Rice, E.Tanke)

  • In parallel to using BPM6W1 for this machine study, also BSM23W was used. Two series of measurements were made: one with current in bunches 3,4 and 5 for trains 1 through 8 (8 by 3), and one with current in bunches in a 6 by 5 arrangement. Data was taken with the bsm_v1 program, option 4 (averaged data), Tglob=1820, Tcard=100, 100 loops and 1000 turn averages. Measurement data are on the server in sub-directories of tanke/cesr/ts101/bsm23w/test_300avg/2006_03_15
  • Measurements in 8 by 3 configuration:
    • /0930 : 0.5 mA per bunch, PMT HV=-440V, seperators on
    • /0939 : 2.0 mA per bunch, PMT HV=-400V, seperators off
    • /0945 : 4.0 mA per bunch, PMT HV=-350V, seperators off
    • /0956 : 0.75 mA per bunch, PMT HV=-440V, seperators off
    • /1006 : 3.0 mA per bunch, PMT HV=-350V, seperators off
  • Measurements in 6 by 5 configuration:
    • /1106 : 1.0 mA per bunch, PMT HV=-400V, seperators off
    • /1111 : 2.0 mA per bunch, PMT HV=-380V, seperators off
    • /1114 : 3.0 mA per bunch, PMT HV=-350V, seperators off
    • /1120 : 4.0 mA per bunch, PMT HV=-330V, seperators off
  • Measurement results are in this file.

March 21, 2006 Geometric channel 20 disfunctional (M.Palmer, E.Tanke)

  • Geometric channel 20 (in 1..32) was found to be disfunctional.

March 22, 2006 Geometric channel 20 functional again (M.Palmer, E.Tanke)

  • Geometric channel 20 (in 1..32) is now functional again, following a re-seating of the PMT array.
  • The PMT rail has be re-aligned such as to give approximately optimal signal for the same mirror settings as for which the light is optimal for the CCD camera.

April 2006

April 24, 2006 (M.Billing, G.Codner, M.Palmer, E.Tanke): Damping time and Ecloud Machine Studies using the BSM at 23W

  • Ecloud measurements with electrons (Morning MS session, under CLEO like conditions)
    • Measurement files are currently stored on CESR29 in the [cesr.palmer.ts101.bsm.saved_data] directory
    • Filenames for electrons are bsm23w_paw_XYZ.dat and bsm23w_cur_XYZ.dat, where XYZ is the year,date and time of the measurement
    • Filenames for positrons are bsm23e_paw_XYZ.dat and bsm23e_cur_XYZ.dat, where XYZ is the year,date and time of the measurement
    • Settings:
      • 10k turns, 100 turns/average, taking B1..B45 and B50 as pedestal
      • HW trigger=0 (no bit set), Tglob=1820, Tcard=100
    • 0.5 mA per bunch, PMT HV=-500V, file 1312 (pinger off)
    • 1.0 mA per bunch, PMT HV=-460V, file 1318 (pinger on)
    • 1.0 mA per bunch, PMT HV=-460V, file 1319 (pinger off)
    • 1.5 mA per bunch, PMT HV=-440V, file 1324 (pinger off)
    • 1.5 mA per bunch, PMT HV=-440V, file 1327 (pinger off, repeat of 1324)

  • Ecloud measurements with electrons (Evening MS session, under CHESS like conditions)
    • Measurement files are currently stored on CESR29 in the [cesr.palmer.ts101.bsm.saved_data] directory
    • Filenames for electrons are bsm23w_paw_XYZ.dat and bsm23w_cur_XYZ.dat, where XYZ is the year,date and time of the measurement
    • Filenames for positrons are bsm23e_paw_XYZ.dat and bsm23e_cur_XYZ.dat, where XYZ is the year,date and time of the measurement
    • Settings:
      • 10k turns, 100 turns/average, taking 6 Trains x 5 Bunches and T6B14 as pedestal
      • HW trigger=0 (no bit set), Tglob=1820, Tcard=100
    • 1.0 mA per bunch, PMT HV=-400V, file 2137
    • 1.0 mA per bunch, PMT HV=-400V, file 2141 (repeat of 2137)
    • 2.0 mA per bunch, PMT HV=-400V, file 2148
    • 4.0 mA per bunch, PMT HV=-380V, file 2152
    • 8.0 mA per bunch, PMT HV=-360V, file 2200

May 2006

May 16, 2006 (M.Palmer et al): Septum effect machine studies

  • 1 bunch of electrons at 1.7 mA, adjusted the timing of the Vert pinger to be roughly at the time of the septum. TIM TRANSFER 21 = 17947. BPM measurements are described ont he BPM6W1 twiki page.
  • PMT HV=-600 V, Tglob=1820, HW trigger=1 (i.e. bit 0 set)
  • VERT PINGER ON cmd = 100
    • BPM Type II multi-file 00323
    • BPM Type II single-file 213
    • BPM Type I 3053
  • Adjusted Coupling SQ4W = -765 --> 546, SQ4E = -1111 --> 289
  • VERT PINGER ON cmd = 100
    • BPM Type II multi-file 00326
    • BSM file 1946:
      • BSM23W_PAW_2006_05_16_1946.GIF:
        BSM23W PAW 2006 05 16 1946.GIF
    • BPM Type II single-file 214
    • BPM Type I 3054
  • E Septum 3 ON cmd = 1355
    • BPM Type II multi-file 00329
    • BSM file 1952:
      • BSM23W_PAW_2006_05_16_1952.GIF:
        BSM23W PAW 2006 05 16 1952.GIF
    • BPM Type II single-file 215
    • BPM Type I 3055
  • Both OFF
    • BPM Type II multi-file 00331
    • BSM file 1959:
      • BSM23W_PAW_2006_05_16_1959.GIF:
        BSM23W PAW 2006 05 16 1959.GIF
    • BPM Type II single-file 216
    • BPM Type I 3056

May 23, 2006 (M.Billing, G.Codner, D.Rice, E.Tanke): Ecloud Machine Study using the BSM at 23W

  • Ecloud related measurements were made with fill 1tx40b or less, e-, Tglob=1820; 10K turns deep with 100 turn averaging (option 6 in VMS BSM_V1 program) as well as 1k turns of raw data (optoion 2), both with 1 added bunch as pedestal); HW trigger=1 (bit 0 set).
  • Measurement files are on the server in directory /nfs/cesr/temp/ts101/BSM23W_data/2006/2006_05
    • 1.50 mA per bunch, PMT HV=-440V, 1tx40b, averaged data file 1900, raw data file 1901, (solenoid off)
    • 0.50 mA per bunch, PMT HV=-490V, 1tx40b, averaged data file 1923, raw data file 1916, (solenoid off)
    • 1.00 mA per bunch, PMT HV=-460V, 1tx40b, averaged data file 1927, raw data file 1928, (solenoid off)

July 2006

July 10, 2006 (E.Tanke): Mirror adjustments following the replacement of the fixed main mirror in the beam pipe

  • Significantly more signal was obtained after re-adjustment of the horizontal mirrors in the periscope:
    MS10Jul2006.jpg

August 2006

August 1, 2006 (R.Holtzapple, E.Tanke): HEP BSM23w Measurements

  • HEP BSM23w Measurements
PMT Voltage    BSM File          Pattern                                Condition
420V              166           23 bunches + 1 pedestal       10,000 turns/100 turn average
420V              167           23 bunches + 1 pedestal       Single turn-500 turns

August 8, 2006 (D.Hartill, M.Palmer, E.Tanke): Installation of new optics line; first measurments

  • A new optics line, including shutter, comparable to the BSM23E was installed at 23W today
  • The DSP was replaced by one having a CM board and ethernet
  • Timing scan was made (file 238) with PMT HV=-450V, timing step size of 20; peak timing now around TglobA=430

August 9, 2006 (E.Tanke): Swap of front-end boards

  • The DSP installed yesterday had been used before for the Xray beam experiments at CHESS hutch B1. It turns out that the front end boards of this DSP have a gain of about a factor 10 lower than the standard BSM boards (the "Xray" boards had been used in the FLM before the FLM boards were swapped for ones that include accumulator boards). Today the front-end boards from the old BSM23W module were installed in the new BSM23W module, thus replacing the low gain for high gain boards. Note:
    • The BSM boards give about 30000 counts (pedestal subtracted) for a 60 mV pulse
    • The old FLM boards give about 3000 counts (pedestal subtracted) for a 60 mV pulse
    • Timing scans for the above 2 cases are shown in this PDF file

August 10, 2006 (E.Tanke): DSP re-installed in tunnel; timing scans

  • The DSP got re-installed in tunnel
  • Timing scans made for T1B3 show that there is a dependency of the optimal timing setting on the PMT voltage (see this PDF file)

August 14, 2006 (E.Tanke): BSM routines now propose timing as a function of selected PMT HV

  • A fit to the BSM23W timing scans of 10-Aug-2006 has been made. The fitted curve (see below) is now used to propose an optimum timing to the user. The timing will change depending on the PMT HV chosen. The same has been done for BSM23E
    timing vs pmthv.gif

August 15, 2006 (E.Tanke): BSM23W Electron beam size calibration

  • Quick scan (from 8:15 am to 8:27 am) with VBUMPING 23 with as purpose calibration of BSM23W, using VSIZE data as reference.
  • BSM related settings : PMT HV=-650V, TglobA=200, 1000 averages of 100 turns
VB23    VSIZE file    BSM23W file
(CU)      number        number
-75       082159        301
-55       082252        302
-35       082352        303
-15       082445        304
+05       082534        305
+25       082633        306
  0       082717        307 (nominal VB23 setting)
  • Calibration results are shown below:
    calibration 15Aug2006.gif
  • Moved measurement files upto and including #307 from the VAX to the server in subdirectories of /nfs/cesr/temp/ts101/BSM23W_data/2006

August 21, 2006 (R.Holtzapple,J.Kern,J.Codner,E.Tanke): MS Beamsize as a function of e- and e+ current

  • Measurements for 10k turns averaged data with 100 turn/avg; also logging of FLM data:
    • Executable: [cesr.flm.data_log]nflm_dlog.exe
    • Input data file: [cesr.flm.data_log]flm_data.inp (or FLM_DATA_2006_08_21.INP)
    • Output files on: log3$disk[cesr.flm.dlog]
    • Enable execution with: fff vxputn FLM FIT DATA 30 30 0
    • Disable execution with: fff vxputn FLM FIT DATA 30 30 1
  • Electrons only:
    • 0.5 mA, FLM logfile #46
      • PMT HV=-650V file #338
      • PMT HV=-700V file #339
      • PMT HV=-750V file #340
    • 1.0 mA, FLM logfile #48
      • PMT HV=-700V file #342
      • PMT HV=-750V file #341
    • 1.5 mA, FLM logfile #49
      • PMT HV=-700V file #343
      • PMT HV=-750V file #344 (some ADC saturation)
    • 2.0 mA, FLM logfile #50
      • PMT HV=-700V file #346
      • PMT HV=-750V file #345 (some ADC saturation)
    • 2.5 mA, FLM logfile #51
      • PMT HV=-700V file #347
      • PMT HV=-680V file #348
  • Electrons and Positrons:
    • 0.5 mA e-,2.5 mA e- , FLM logfile #58
      • PMT HV=-700V file #361
    • 1.0 mA e-,2.5 mA e- , FLM logfile #59
      • PMT HV=-700V file #364
    • 1.5 mA e-,2.5 mA e- , FLM logfile #60
      • PMT HV=-700V file #365
    • 2.9 mA e-,2.5 mA e- , FLM logfile #61
      • PMT HV=-650V file #370

August 24, 2006 (G.Codner,M.Forster,M.Palmer,E.Tanke) Beam splitters and MS ECLOUD and wiggler

  • Beam splitter work:
    • Inspected the first beam splitter (CCD line), which was found to be a 50T/50R; left it untouched
    • Replaced the second beam splitter (PMT line), which was a 92T/8R; replaced it by a 50T/50R
    • Note that the shutter is followed by a 500nm filter and an expander lens on seperate posts
    • File 400: 1k turns of raw data, PMT HV=550V, 9x5 with 9 pedestals (but beam was 1 train of 45 bunches); analysis of this file shows that we now have sufficient light on the PMT for good turn-by-turn data ( RMS less than 10% of pulse height) !
  • MS ECLOUD
    • 1.5 mA, 45 bunches and 1 pedestal, PMT HV=550V, T=210
      • File 401: 1k turns of raw data
      • File 402: 10k turns of averaged data, 100 turns/avg, 100 averages
    • 0.5 mA, 45 bunches and 1 pedestal, PMT HV=550V, T=210
      • File 404: 1k turns of raw data
      • File 405: 10k turns of averaged data, 100 turns/avg, 100 averages
      • File 406: timing scan at PMT=600V
      • File 407: 1k turns of raw data with PMT HV=600V
  • MS Wiggler test, 0.5 mA, 45 bunches and 1 pedestal, PMT HV=600V, T=200, 6 wigglers off
    • File 429: 10k turns of averaged data, 100 turns/avg, 100 averages
    • File 430: 1k turns of raw data

August 31, 2006 (R.Holtzapple,J.Kern,E.Tanke) BSM measurements during CLEO conditions (3686 run)

  • Parasitic CESR-c Colliding Beams PMT Measurements; 24 bunches HEP, Bunch Pattern ~8x3:
    • Bunch 3 Trains 1-8
    • Bunch 4 Trains 1-9
    • Bunch 5 Trains 2-8
  • Measurements made both for electrons and positrons
Condition PMT Turns Data Shutter delay
8x3 e+ 100avg 550 10000 451 50K turns
8x3 e- 100avg 550 10000 452 50K turns
8x3 e+ Raw 550 500 454 50K turns
8x3 e- Raw 550 500 453 50K turns
8x3 e+ Raw 550 10000 456 50K turns
8x3 e- Raw 550 10000 455 50K turns
T3 B3 e- Raw 550 20000 458 60K turns
8x3 e+ Raw 550 10000 460 60K turns
8x3 e- Raw 550 10000 461 60K turns
Train4 b3-5 e+ Raw 550 100000 463 70K turns
Train4 b3-5 e- Raw 550 100000 462 70K turns
  • Files up to and including file number 449 are now on the server ( /nfs/cesr/temp/ts101/BSM23W_data )

September 2006

September 6, 2006 (M.Palmer, E.Tanke): Modifications to BSM23W optics line

  • Replaced the expander lenses in both the east and west PMT optics lines to reduce the magnification by approximately two-thirds f=-25 mm lens replaced with an f=-50 mm lens in each line but with longer moment arm) . In addition replaced the two PMT optics line mirrors in the east with larger aperture mirrors (old mirrors had been an operational limit in the east). Optics alignment was maintained using the laser alignment setup so we should be ready to go this evening.

September 7, 2006 (M.Palmer, L.Schachter, R.Holtzapple, G.Codner, E.Tanke): ECLOUD measurements

  • Measurement types
    • (a): 100 turn avg, 10k turns total, 45 bunches + 1 pedestal
    • (b): 1k turns of raw data, 45 bunches + 1 pedestal
    • (c): 10k turns of raw data, selected bunches + 1 pedestal
  • Electron measurements 1.25 mA , PMT HV=550V, T=230, selected wigglers off, feedback on
    • (a): 527
    • (b): 528
    • (c): 530 (bunches 1,11,21,27-31,34,37 +1 pedestal)
  • Electron measurements 0.25 mA , PMT HV=600V, T=200, selected wigglers off, feedback on
    • (a): 531
    • (b): 532
    • (c): 533 (bunches 1,11,21,27-31,34,37 +1 pedestal)
  • Electron measurements 0.75 mA , PMT HV=600V, T=200, selected wigglers off, feedback on
    • (a): 534
    • (b): 535
    • (c): 536 (45 bunches +1 pedestal)
  • Electron measurements 1.25 mA , PMT HV=550V, T=230, all wigglers on, feedback on
    • (a): 554
    • (b): 555
  • Electron measurements 0.25 mA , PMT HV=600V, T=200, all wigglers on, feedback off
    • (a): 571
    • (b): 572
  • Electron measurements 0.25 mA , PMT HV=600V, T=200, all wigglers on, feedback on
    • (a): 573
    • (b): 574
    • (c): 575 (bunches 1,6,10,11,12,13,14,17,20 +1 pedestal)
  • Electron measurements 0.75 mA , PMT HV=600V, T=200, all wigglers on, feedback off
    • (a): 576
  • Electron measurements 0.75 mA , PMT HV=600V, T=200, all wigglers on, feedback on
    • (a): 577
    • (c): 578 (45 bunches +1 pedestal)
  • Files up to and including file number 519 are now on the server ( /nfs/cesr/temp/ts101/BSM23W_data )

September 18, 2006 (E.Tanke) Calibration scan

  • Calibration scan with VB23 with CCD and BSM data taking; 2 mA, T1B1 (single bunch)
    • VB23=+5, CCD file 081440, BSM file 617
    • VB23=-35, CCD file 081609, BSM file 618
    • VB23=-15, CCD file 081655, BSM file 619
    • VB23=+5, CCD file 081728, BSM file 620
    • VB23=+25, CCD file 081802, BSM file 621
    • VB23=+45, CCD file 081834, BSM file 622
    • VB23=+65, CCD file 081931, BSM file 623
    • Note that VB23=+5 was the nominal setting
  • Based on the positional change, the magnification factor is 8.6
  • Based on the beamsize, the magnification factor is 6.9
  • Files up to and including file number 623 are now on the server ( /nfs/cesr/temp/ts101/BSM23W_data )

September 26, 2006 (M.Palmer, L.Schachter, G.Codner, M.Pivi, D.Rice, E.Tanke): ECLOUD measurements at CHESS energy

  • Measurement types
    • (a): 100 turn avg, 10k turns total, 45 bunches + 1 pedestal
    • (b): 1k turns of raw data, 45 bunches + 1 pedestal
    • (c): 10k turns of raw data, 45 bunches + 1 pedestal
  • Electron measurements 0.5 mA , PMT HV=500V, T=300
    • (a): 664
    • (b): 665
  • Electron measurements 0.85 mA , PMT HV=500V, T=300
    • (a): 666
    • (b): 667
      • Totals (mA) (e- and e+): 32.88000 0.7050000
      • Beamsizes (um) (e- and e+): 249.0000 7.000000
      • SYLTOUCHY E- HORZNTAL TILT= 683
      • SYLTOUCHY E- VERTICAL TILT= 489
      • SYLTE CON E- HORZNTAL TILT= 540
      • SYLTE CON E- VERTICAL TILT= 587
      • SYLTE CON E- FOCUSING = 415
      • SYLTE CON E- INTERFEROMETER= 453
      • E- HORZ FEEDBK = -2047
      • E- VERT FEEDBK = -1100
    • (c): 668 Taken after refill; chromaticity (CSR XQUNEING 1 1) was changed from 972 to 520
      • Totals (mA) (e- and e+): 31.21600 0.6530000
      • Beamsizes (um) (e- and e+): 105.0000 7.000000
      • SYLTOUCHY E- HORZNTAL TILT= 682
      • SYLTOUCHY E- VERTICAL TILT= 489
      • SYLTE CON E- HORZNTAL TILT= 540
      • SYLTE CON E- VERTICAL TILT= 587
      • SYLTE CON E- FOCUSING = 415
      • SYLTE CON E- INTERFEROMETER= 453

October 2006

October 02, 2006 (E.Tanke): New BSM_V1 executable in DEV directory

  • There is now a new bsm_v1 executable, which generates bsmYYY_cur_xxxx.dat files that contain more info about the machine status and saveset numbers.
  • This executable also has linked to it a first generation CBI_NET software, allowing for ethernet transfer (for "raw" data only)
  • Note that this executable is located in the [cesr.palmer.ts101.bsm.dev] directory

October 03, 2006 (M.Forster, R.Holtzapple, J.Kern, E.Tanke): Beamsize measurements at CHESS energy under different coupling conditions

  • Measurement types
    • (a): 100 turn avg, 10k turns total
    • (b): 1k turns of raw data
  • Filled bunches: T1B1..B5 -> T4B1..B5 positrons 7 mA/bunch; T5B1..B5 -> T9B1..B5 electrons 7 mA/bunch; Electron measurements:
    • (a): 707
    • (b): 709
  • Change coupling; Electron measurements:
    • (a): 711
    • (b): 713
  • Electron bunches: T5B1..B5 -> T9B1..B5 electrons 3.5 mA/bunch (also positrons 4x5 in CESR); Electron measurements:
    • (a): 721
    • (b): 723
  • Change coupling; Electron measurements:
    • (a): 725
    • (b): 727
  • Only keep electron bunches: T8B1..B5 -> T9B1..B5 electrons 9 mA/bunch (positrons 4x5 still in CESR); Electron measurements:
    • (a): 729
    • (b): 731
  • Change coupling; Electron measurements:
    • (a): 733
    • (b): 735

October 10, 2006 (M.Forster, R.Holtzapple, J.Kern, E.Tanke): Beamsize measurements at CHESS energy under different coupling conditions

  • Measurement types
    • (a): 100 turn avg, 10k turns total
    • (b): 1k turns of raw data
  • First measurements, 9x5, positrons only, 4.75 mA/bunch
  • Adding electrons 9x5 1.5 mA/bunch; Electron measurements:
    • (a): 759
    • (b): 761
  • Change to poorly corrected coupling; Electron measurements:
    • (a): 763
    • (b): 765
  • Change to well corrected coupling; Electron measurements:
    • (a): 767
    • (b): 769
  • Increase electron bunch current to 3 mA/bunch; Electron measurements:
    • (a): 771
    • (b): 773
  • Change to poorly corrected coupling; Electron measurements:
    • (a): 775
    • (b): 777
  • Change to well corrected coupling; Electron measurements:
    • (a): 779
    • (b): 781
  • Increase electron bunch current to 4.75 mA/bunch; Electron measurements:
    • (a): 783
    • (b): 785
  • Electrons are falling out. Refill electrons (and positrons); Electron measurements:
    • (a): 787
    • (b): 789
  • Change to poorly corrected coupling; Electron measurements:
    • (a): 791
    • (b): 793
  • Change to well corrected coupling; Electron measurements:
    • (a): 795
    • (b): 797
  • Files up to and including file number 749 are now on the server ( /nfs/cesr/temp/ts101/BSM23W_data )

October 12, 2006 (R.Holtzapple, J.Kern, E.Tanke): Beamsize measurements at CHESS energy

  • Measurement types
    • (a): 100 turn avg, 10k turns total
    • (b): 1k turns of raw data
    • (c): 10k turns of raw data
  • First measurements, positrons 9x5 and electrons 9x5
  • Electron measurements:
    • (a): 799
    • (b): 801
  • Adjust SYLTE mirrors; Electron measurements:
    • (a): 803
    • (b): 805
    • (c): 808
  • After tests conducted by JPS (which involved moving the mirrors), mirrors set back to initial set values; Electron measurement:
    • (c): 811
  • Files up to and including file number 811 are now on the server ( /nfs/cesr/temp/ts101/BSM23W_data )

October 17, 2006 (J.Codner,M.Forster, R.Holtzapple, J.Kern, E.Tanke): Beamsize measurements at CHESS energy under different coupling conditions

  • Measurement types
    • (a): 100 turn avg, 10k turns total
    • (b): 1k turns of raw data
  • First measurements after down recovery, 9x5, positrons and electrons, 4.75 mA/bunch ; electron measurements:
    • (a): 818
    • (b): 822
  • Set initial settings of 10-Oct-2006 (saveset 112793); change settings (store new saveset) and measure with electrons only in CESR * (a): 824 * (b): 825
  • Change to poorly corrected coupling; Electron measurements:
    • (a): 826
    • (b): 827
  • Change to well corrected coupling; Electron measurements:
    • (a): 828
    • (b): 829
  • Reset to initial settings with electrons and positrons; Electron measurements: * (a): 830 * (b): 832
  • Change to well corrected coupling; Electron measurements:
    • (a): 834
    • (b): 836
  • Change to poorly corrected coupling; Electron measurements:
    • (a): 838
    • (b): 840

October 24, 2006 (M.Forster, R.Holtzapple, J.Kern, E.Tanke): Beamsize measurements at CHESS energy under different coupling conditions

  • Measurement types
    • (a): 100 turn avg, 10k turns total
    • (b): 1k turns of raw data
  • First measurements after down recovery, 9x6, positrons and electrons, 3.8 mA/bunch ; electron measurements:
    • (a): 851
    • (b): 853
  • Measurements with positrons only in CESR
  • Refill and measure with positrons AND electrons in CESR; Electron measurements:
    • (a): 864
    • (b): 868
  • Change to poorly corrected coupling; Electron measurements:
    • (a): 869
    • (b): 871
  • Change to well corrected coupling; Electron measurements:
    • (a): 873
    • (b): 875
  • Both RF phases changed by +4 cu; Electron measurements:
    • (a): 877
    • (b): 879

October 31, 2006 (G.Codner, M.Forster, R.Holtzapple, J.Kern, E.Tanke): Beamsize measurements at CHESS energy under different coupling conditions, V tune change

  • Measurement types
    • (a): 100 turn avg, 10k turns total
    • (b): 1k turns of raw data
  • First measurements after down recovery, 9x6, positrons and electrons, 4 mA/bunch ; electron measurements:
    • (a): 888
    • (b): 890
  • Change to mangled coupling; electron measurements:
    • (a): 892
    • (b): 894
  • Change to well corrected coupling; electron measurements:
    • (a): 896
    • (b): 898
  • Change V tune to 1 kHz below nominal ; electron measurements:
    • (a): 900
    • (b): 902
  • Change V tune to 1 kHz above nominal ; electron measurements:
    • (a): 904
    • (b): 906
  • Change V tune to 2 kHz below nominal ; electron measurements:
    • (a): 908
    • (b): 910

Topic revision: r45 - 21 May 2007, EugeneTanke
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding CLASSE Wiki? Send feedback